OpenStudy (anonymous):

Angle A is 21 less than twice as large as its supplement, Angle B. Find the measure of each angle.

OpenStudy (anonymous):

PLEASE HELP ME!! FOR FAN AND MEDAL

OpenStudy (stefrheart):

First you need to try to understand what they are saying. Start with What is a supplementary angle?

OpenStudy (anonymous):

A supplementary angle is an angle that adds up to 180 degrees

OpenStudy (stefrheart):

Yes so you know that Angle A and Angle B is equal to 180 degrees right? Now you need to make an equation 2x-21+x=180

OpenStudy (anonymous):

Okay so I think I understand it from here. 2x-21+x=180 +21 +21 2x+x=180 3x=180 ------- 3 x=60

OpenStudy (stefrheart):

You are on the right track but you forgot to add the 21 so it would look like this 2x-21+x=180 +21 + 21 2x+x=201 3x=201 x=67

OpenStudy (stefrheart):

Now remember we were solving x and we are not done yet

OpenStudy (anonymous):

Sorry I'm tired. Haha. I know that's the right process. I just overlooked it. THanks for pointing it out though

OpenStudy (stefrheart):

Lol its okay :) i know how you feel im just getting over the flu!

OpenStudy (stefrheart):

So now to finish it its Angle A= 2x-21 Angle B= x

OpenStudy (stefrheart):

Then you are done! One problem that is over with!

OpenStudy (anonymous):

Thanks! I have a couple more. Do you think you might be able to help?

OpenStudy (stefrheart):

Absolutely! and Welcome!:)

OpenStudy (anonymous):

Gimme one second and I'll figure out which ones I need help with

OpenStudy (stefrheart):

Lol ill be here probably!

OpenStudy (anonymous):

Angle A is congruent to its supplement, Angle B.

OpenStudy (stefrheart):

okay is that the whole question?

OpenStudy (anonymous):

Yes

OpenStudy (stefrheart):

Oh okay so since its congruent its the same right? So it would just be \[180\div2\]

OpenStudy (anonymous):

So 90? seems too easy. Lol

OpenStudy (stefrheart):

lol thats what i thought! Thats why i asked if that was the whole question

OpenStudy (anonymous):

Thats all thats there. :) Haha. The next question is Angle A is 18 more than five times its complement Angle B

OpenStudy (stefrheart):

Hahahaha! Okay back to buisness... you know what a complement is right?

OpenStudy (anonymous):

90 degree measurement?

OpenStudy (stefrheart):

yes its two angles that add up to 90 degrees! so 5x+18+x=90

OpenStudy (anonymous):

So x=12?

OpenStudy (stefrheart):

12? i got 18.. 5x+18+x=90 6x+18=90 6x=108 x=18

OpenStudy (stefrheart):

Then to finish it: Angle A= 5x+18 Angle B= x

OpenStudy (anonymous):

But since in the original equation 18 is positive, wouldn't you subtract it to balance both sides?

OpenStudy (stefrheart):

Omg! i feel so stupid yes you are correct! Im so sorry!! Ahh!!

OpenStudy (anonymous):

No problem. I've made the same mistake before. :) I appreciate your help. :)

OpenStudy (stefrheart):

Lol welcome! Even though i messed up!

OpenStudy (stefrheart):

Is there any more?

OpenStudy (anonymous):

No not on this assignment. But what would the equation be now that I know x=12?

OpenStudy (stefrheart):

Angle A should be 78 degrees and Angle B should be 12 degrees

OpenStudy (anonymous):

Thanks!

OpenStudy (stefrheart):

Welcome! Im happy to help! :)